### v1

@@ 1 Introduction
Section 2 introduces our proposed voting model and relevant axiomatic properties. Section 3 presents an exhaustive analysis of the different variations of our central axiom — anchor-proofness — and demonstrates impossibility and possibility results for each one of them. Section  4 presents the model where a social planner can manipulate, and provides certain characterization results. Section 5 concludes and discusses future lines of research.
@@ 4 Strategies for the Social Planner
We then relax the assumption of full information and study precisely how the planner’s ability to manipulate the outcome depends on the information she possesses about voters’ preferences. We first show that, under zero information, manipulation is impossible for any rule satisfying weak and total unanimity, SAV being one of them (Theorem 7), and also for the nomination rule (Theorem 10). We then demonstrate that even limited aggregate information — such as acceptability points or plurality information — can restore manipulability for broad classes of rules (Theorems 8911, and  12).
@@ 4.1 Anchor-proof Profiles under Full Information
It should be noted that when adding total unanimity to the requirements of Theorem 6, the result remains the same. This is because total unanimity only plays a role in the particular profile where every approval ballot is exactly $X$. The profiles employed in the characterisation already omit this profile.
@@ 4.2 Anchor-proofness under Partial Information
So far, we have been able to find a small number of profiles for which particular rules are anchor-proof, suggesting that voters wield only limited influence. Crucially however, these results rest on a strong assumption about the information available to the planner: anchor-proofness fails precisely because the planner might be able to select a specific presentation order relying on the intrinsic preferences of the voters.
Next, we turn to settings where the planner might not have full information about voters’ preferences. Using the model of Reijngoud and Endriss (2012), we define a notion called ‘information function’ that maps each preference profile $\boldsymbol{p}$ to a type of information. Specific such functions $\mathit{info}$ that we will consider are the following:
@@
We ask whether, under partial information, the social planner can manipulate the voting outcome to her advantage. Given a set of possible profiles $W\subseteq P(X)^{N}$ consistent with the information of the social planner and a voting rule, we call a strategy optimal for the social planner if $(i)$ for every possible profile in $W$ it induces an outcome that is at least as good as the outcome of every other strategy, and $(ii)$ in some possible profile in $W$ it induces an outcome that is strictly better than the outcome of some other strategy. [^4] Case $(ii)$ is important to avoid the scenario where all strategies give the same outcome in all possible profiles, meaning that none of them can be considered optimal. If the social planner has an optimal strategy given her information, then she can manipulate the voting rule. The following definition formalizes this idea.
@@
Note that except for the case of zero information, all information functions described in this section allow us to distinguish the status of the different alternatives in a profile (which is sensible in several contexts of collective decisions, including polls of political elections and preliminary results of participatory budgeting contests). But in other settings, we can imagine that the provided information does not allow the planner to separate the different alternatives. The reader might suspect that for reasonable rules such as SAV, manipulability is then avoidable, since the planner has to fix a presentation order of the alternatives in her strategy without being able to exclude any permutation of the alternatives in the profiles she considers possible. However, this conjecture does not hold in general as Example 7 shows.
@@ 5 Conclusion
Our analysis of anchor-proof voting rules demonstrates how axiomatic characterizations in social choice can be extended to accommodate systematic deviations from rational preference formation. We have started by considering voters with intrinsic preference rankings who are subject to an anchoring-bias, influencing the formation of their approval ballots via the presentation order of the alternatives. Our central characterization result shows that no approval-based voting rule (except constant ones) is robust to the impact of such an anchoring-bias. On the positive side, we also find that a social planner who has no information about the voters’ underlying preferences cannot choose the presentation order to safely steer the voting outcome to her advantage.
A natural extension of our framework is to consider ballots with richer structures, particularly ranked ballots. In such settings, the derived ballots under an anchoring bias would typically correspond to incomplete preference orders over the set of alternatives. A preliminary conjecture is that every tops-only rule — that only takes into account the most preferred alternatives of each voter to determine the outcome — is anchor-proof, since the most preferred alternative always appears in the corresponding derived ballot regardless of the presentation order. This conjecture naturally connects the present work to the broader literature on collective decision-making under incomplete preferences, where the interplay between informational incompleteness and behavioral robustness may yield novel characterization results.
A further direction involves exploring alternative mechanisms for deriving anchor-biased ballots. One possibility is to consider trichotomous ballots, where alternatives are classified as approved, disapproved, or neutral. Under this formulation, an alternative that is accepted by a voter but appears after a more preferred one could be treated as neutral rather than disapproved, offering a more fine-grained representation of cognitive constraints.
@@
The negative results established in our paper motivate yet another research direction: designing voting rules with quantitatively optimal degrees of immunity to specific biases. From a computational social choice perspective, understanding which biases can be efficiently detected or corrected through algorithmic pre-processing of preference data, versus which require fundamentally different aggregation mechanisms, has important practical implications for both human and multi-agent collective-decision contexts.

### v2

@@ 1 Introduction
Section 2 introduces our proposed voting model and relevant axiomatic properties. Section 3 presents an exhaustive analysis of the different variations of our central axiom — anchor-proofness — and demonstrates impossibility and possibility results for each one of them. Section  4 presents the model where a social planner can manipulate, and provides certain characterization results. Section  5 discusses some extensions of the model and open questions. Section 6 concludes and discusses future lines of research.
@@ 4 Strategies for the Social Planner
We then relax the assumption of full information and study precisely how the planner’s ability to manipulate the outcome depends on the information she has on the preferences of the voters. We first show that, under zero information, manipulation is impossible for any rule satisfying weak and total unanimity, SAV being one of them (Theorem 7), and also for the nomination rule (Theorem 10). We then demonstrate that even limited aggregate information — such as acceptability points or plurality information — can restore manipulability for broad classes of rules (Theorems 8911, and  12).
@@ 4.1 Anchor-proof Profiles under Full Information
It should be noted that when adding total unanimity to the requirements of Theorem 6, the result remains the same. This is because total unanimity only plays a role in the particular profile where every approval ballot is exactly $X$. The profiles employed in the characterization already omit this profile.
@@ 4.2 Anchor-proofness under Partial Information
So far we have been able to find only a small number of profiles for which particular rules are anchor-proof. Crucially, these results rest on a strong assumption about the information available to the planner: anchor-proofness often fails precisely because the planner might be able to select a specific presentation order relying on the intrinsic preferences of the voters.
Next, we turn to settings where the planner lacks full information about the voters’ preferences. Using the model of Reijngoud and Endriss (2012), we define a notion called ‘information function’ $\mathit{info}$ that maps each preference profile $\boldsymbol{p}$ to a type of information. We will study the following such functions:
@@
We ask whether, under partial information, the social planner can manipulate the voting outcome to her advantage. Given a set of possible profiles $W\subseteq P(X)^{N}$ consistent with the information of the social planner and a voting rule, we call a strategy optimal for the social planner if $(i)$ for every possible profile in $W$ it induces an outcome that is at least as good as the outcome of every other strategy, and $(ii)$ in some possible profile in $W$ it induces an outcome that is strictly better than the outcome of some other strategy. [^4] Case $(ii)$ is important to avoid the scenario where all strategies give the same outcome in all possible profiles, meaning that none of them can be considered optimal. If the social planner has an optimal strategy given her information, then she can manipulate the voting rule. Definition 8 formalizes this idea.
@@
5 Discussion
In this section, we discuss two extensions of our framework. First, we examine the setting with ranked ballots rather than approval ballots. Second, we analyze a particular class of information functions in which alternatives are indistinguishable to the social planner.
5.1 Ranked Ballots
A natural variation of our setting assumes that voters generate top-truncated ranked ballots rather than approval ballots. A top-truncated ranking constitutes a linear order over some subset of the alternatives, with the implicit assumption that all non-ranked alternatives are inferior to all ranked alternatives. [^5] Top-truncated rankings are commonly studied in the social choice literature about incomplete preferences (Baumeister et al., 2012; Terzopoulou and Endriss, 2021). The anchoring process here operates similarly: given an intrinsic preference of a voter that is a complete ranking (possibly endowed with an acceptance threshold), when a presented alternative is less preferred than some alternative already ranked in the voter’s ballot, then it is discarded; otherwise, if it is accepted, it is ranked in accordance with the voter’s intrinsic preference.
To be more precise, a few definitions are in order. The top-truncated ballot of voter $i$ with intrinsic preference $p_{i}$ given the presentation order $\sigma_{i}$ is denoted $p_{i,\sigma_{i}}$ (a linear order over a subset of $X$). For a ranking $p$ (truncated or complete), $p^{1}$ is its top alternative, meaning the most preferred alternative in that ranking.
We introduce the notion of a rank-based voting rule $F_{r}$, which takes as input a profile of (possibly truncated) ballots and outputs a subset of winning alternatives. We say that $F_{r}$ satisfies tops-only if for every two profiles $\boldsymbol{p}$ and $\boldsymbol{q}$ such that $p^{1}_{i}=q^{1}_{i}$ for all $i\in N$, it holds that $F_{r}(\boldsymbol{p})=F_{r}(\boldsymbol{q})$. For a tops-only rule, the only relevant information lies at the top of each voter’s ranking. Thus, it is a rule that does not demand the voters too many details about their preferences. [^6] For further motivation regarding this property, we refer the reader to Sekiguchi (2012). The most famous tops-only voting rule is plurality, which selects as winners the alternatives that occupy most often the top positions in the voters’ preferences of a given profile.
Our main result characterizes all anchor-proof ranked-based rules. Recall that the presentation order $\boldsymbol{\bar{\sigma}}$ shows each voter their top alternative first.
Theorem 13.
A rank-based voting rule is anchor-proof if and only if it is tops-only.
Proof.
(if) Let $F_{r}$ be a tops-only rule. By the natural extension of Remark 1 to top-truncated ballots, we have that $p^{1}=p^{1}_{\sigma}$ for every presentation order $\sigma$. In particular, this holds for every individual preference $p$. Since the rule is tops-only, we then have that $F_{r}(\boldsymbol{p_{\sigma}})=F_{r}(\boldsymbol{p_{\sigma^{\prime}}})$ for all pairs of presentation orders $\boldsymbol{\sigma}$ and $\boldsymbol{\sigma^{\prime}}$. Thus, $F_{r}$ is anchor-proof.
(only if) Let $F_{r}$ be anchor-proof. We need to show that for any two profiles $\boldsymbol{p}$ and $\boldsymbol{q}$ such that $p^{1}_{i}=q^{1}_{i}$ for all $i\in N$, it holds that $F_{r}(\boldsymbol{p})=F_{r}(\boldsymbol{q})$. By anchor-proofness, we have that $F_{r}(\boldsymbol{p})=F_{r}(\boldsymbol{p_{\bar{\sigma}}})$ and $F_{r}(\boldsymbol{q})=F_{r}(\boldsymbol{q_{\bar{\sigma}}})$. Since $\boldsymbol{p_{\bar{\sigma}}}=\boldsymbol{q_{\bar{\sigma}}}$, it follows that $F_{r}(\boldsymbol{p})=F_{r}(\boldsymbol{q})$. We conclude that $F_{r}$ satisfies tops-only. ∎
At first glance, one might expect that finding anchor-proof rules for ranked ballots would be even more challenging than for approval ballots, given the difficulties we encountered in the approval setting. Theorem 13 reveals the contrary: Under ranked ballots, we can construct anchor-proof voting rules by paying attention to those that use only limited information — it is for such informational limitations that the plurality rule has been largely criticized (Laslier, 2011), and it is precisely for those that it is deemed desirable with respect to anchor-proofness. Given a tops-only rule, even a fully informed social planner cannot manipulate the outcome. Thus, an informational trade-off emerges: rules that demand more information provide the social planner with greater leverage for manipulation, while informationally austere rules prove by construction more robust to strategic presentation orders.
5.2 Alternative-Ignorant Information Functions
@@
Note that except for the case of zero information, all information functions described in this Section 4.2 allow us to distinguish the status of the different alternatives in a profile (which is sensible in several contexts of collective decisions, including polls of political elections and preliminary results of participatory budgeting contests). But in other settings, we can imagine that the provided information does not allow the planner to separate the different alternatives. The reader might suspect that for reasonable rules such as SAV, manipulability is then avoidable, since the planner has to fix a presentation order of the alternatives in her strategy without being able to exclude any permutation of the alternatives in the profiles she considers possible. However, this conjecture does not hold in general as Example 7 shows.
@@ 6 Conclusion
Our analysis of anchor-proof voting rules demonstrates how axiomatic characterizations in social choice can be extended to accommodate systematic deviations from rational preference formation. We have started by considering voters with intrinsic preference rankings who are subject to an anchoring bias, influencing the formation of their approval ballots via the presentation order of the alternatives. Our central characterization result shows that no approval-based voting rule (except constant ones) is robust to the impact of such an anchoring bias. On the positive side, we also find that a social planner who has no information about the voters’ underlying preferences cannot choose the presentation order to safely steer the voting outcome to her advantage.
We have further shown that the picture changes substantially when considering ballots with richer structures, particularly ranked ballots. In such settings, the derived ballots under an anchoring bias correspond to incomplete preference orders over the set of alternatives. Our characterization establishes that a rank-based voting rule is anchor-proof if and only if it is tops-only—that is, if it determines the outcome solely based on each voter’s most preferred alternative. This result reveals an informational trade-off: rules that demand more information provide the social planner with greater leverage for manipulation, while informationally austere rules prove inherently more robust to strategic presentation orders. The plurality rule, as a tops-only rule, is thus anchor-proof in the ranked ballot setting, connecting our work to the broader literature on collective decision-making under incomplete preferences.
A future research direction involves exploring alternative mechanisms for deriving anchor-biased ballots. One possibility is to consider trichotomous ballots, where alternatives are classified as approved, disapproved, or neutral. Under this formulation, an alternative that is accepted by a voter but appears after a more preferred one could be treated as neutral rather than disapproved, offering a more fine-grained representation of cognitive constraints.
@@
The negative results established in our paper motivate yet another research direction: designing voting rules with quantitatively optimal degrees of immunity to specific biases. From a computational social choice perspective, understanding which biases can be efficiently detected or corrected through algorithmic pre-processing of preference data, versus which require fundamentally different aggregation mechanisms, has important practical implications for both human and multi-agent collective-decision contexts. This direction can also be considered for ranked ballots, as even though we show that the family of tops-only rules is anchor-proof, one could be interested in optimal informational trade-offs for more complex voting rules.